### v1

# ORACLE-Grasp: Zero-Shot Task-Oriented Robotic Grasping using Large Multimodal Models

Abstract
Grasping unknown objects in unstructured environments remains a fundamental challenge in robotics, requiring both semantic understanding and spatial reasoning. Existing methods often rely on dense training datasets or explicit geometric modeling, limiting their scalability to real-world tasks. Recent advances in Large Multimodal Models (LMMs) offer new possibilities for integrating vision and language understanding, but their application to autonomous robotic grasping remains largely unexplored. We present ORACLE-Grasp, a zero-shot framework that leverages LMMs as semantic oracles to guide grasp selection without requiring additional training or human input. The system formulates grasp prediction as a structured, iterative decision process, using dual-prompt tool calling to first extract high-level object context and then select task-relevant grasp regions. By discretizing the image space and reasoning over candidate areas, ORACLE-Grasp mitigates the spatial imprecision common in LMMs and produces human-like, task-driven grasp suggestions. Early stopping and depth-based refinement steps further enhance efficiency and physical grasp reliability. Experiments demonstrate that the predicted grasps achieve low positional and orientation errors relative to human-annotated ground truth and lead to high success rates in real-world pick up tasks. These results highlight the potential of combining language-driven reasoning with lightweight vision techniques to enable robust, autonomous grasping without task-specific datasets or retraining.
I Introduction
Robotic grasping remains a core challenge in autonomous manipulation, especially in unstructured and novel environments. Traditional approaches often depend on precise object models, depth-based geometry recognition, or supervised learning pipelines trained on large-scale datasets [1, 2]. However, real-world deployment frequently requires interacting with novel objects exhibiting unpredictable shapes with task-based reasoning. In such scenarios, accurate grasp localization, identifying grasp points and orientations that are both feasible and task-appropriate, becomes critical [3]. While recent deep learning methods have advanced grasp quality prediction [4], they primarily focus on geometric properties and neglect semantic reasoning about the object or intended task. This gap motivates the development of new frameworks that combine high-level semantic understanding with spatial reasoning, enabling more human-like, adaptive grasp selection in the real-world.
Figure 1: Overview of the ORACLE-Grasp framework. An RGB-D camera captures the scene. An LMM uses dual prompts over a grid to identify a graspable region, which is then refined with depth data for precise grasp prediction and executed by a robot.
Previous approaches have typically relied on geometric analysis of depth data [5, 6], model-based grasp synthesis [7], or extensive training of deep networks on large-scale grasp datasets [8, 9]. Early work on learning-based methods for grasping novel objects, used machine learning techniques to predict effective grasp points directly from 2D images without requiring complete 3D reconstruction or object recognition [10]. However, deep learning-based approaches typically focus on learning and predicting the quality of candidate grasps, without integrating semantic reasoning about the task or the object’s context [11]. While recent methods have made significant advances in pixel-wise grasp prediction and affordance learning [12, 13], they often lack the semantic reasoning needed for task-oriented grasping, especially in zero-shot settings where the object has not been seen before.
Large Language Models (LLMs) have significantly advanced robotics by enabling natural language understanding for task planning, human-robot interaction, and high-level decision-making [14, 15]. These models allow robots to interpret abstract instructions (e.g., ”organize the desk”) and generate structured action sequences with explainable reasoning. Building on these capabilities, recent developments have introduced Large Multimodal Models (LMMs), which integrate vision and language processing to tackle embodied tasks [16, 17]. By jointly reasoning over visual inputs and textual prompts, LMMs extend the functionality of LLMs, enabling robots to perceive, interpret, and act within their environments with greater semantic understanding.
Prominent examples of LMMs include CLIP [18], Flamingo [19], and LLaVA [20], which have demonstrated impressive capabilities in grounding visual scenes through natural language reasoning [21]. These models are typically trained on large-scale datasets comprising paired image and text data, allowing them to align visual and linguistic representations within a shared semantic space. As a result, LMMs can generalize across domains, leverage contextual information, and infer high-level semantic relationships that are critical for flexible and task-driven robotic behaviors [22]. Recently, LMMs have begun to be applied in robotics to bridge perception, reasoning, and action [16, 23, 24], enabling new capabilities such as instruction following, semantic scene understanding, and object-centric manipulation in real-world environments [25, 26]. By leveraging their ability to reason jointly over visual and linguistic modalities, LMMs offer a promising path toward more adaptable, scalable, and intuitive robotic systems that can operate with minimal task-specific training and additional data collection.
Despite notable advances in high-level robotic task planning and embodied reasoning, the application of LMMs to robotic grasping remains relatively underexplored. Most existing work leveraging LMMs in robotics focuses on semantic instruction following, policy learning, or general manipulation planning, rather than on fine-grained grasp localization and selection. Among the few exceptions, PhyGrasp [27] introduces an LMM that integrates 3D point clouds and language to inform grasping decisions based on estimated physical properties such as material, fragility, and mass. However, PhyGrasp requires detailed 3D shape modeling and the availability of physical property annotations. It also depends on large-scale data collection and does not explicitly address task-driven, human-like grasp selection from single-view visual inputs. Another emerging line of work employs LMMs to direct grasping based on human verbal instructions, providing external context and object knowledge to guide the robot’s actions [28, 29, 30, 31]. While these approaches support task-specific grasping, they depend on human guidance at inference time and often require pretraining on large, specialized grasp datasets, limiting their scalability for fully autonomous deployment. Moreover, they focus on selecting which object to grasp, rather than identifying task-driven, human-like grasp locations on the object itself.
In this letter, we explore the capability of LMMs to infer feasible, task-relevant grasps from single images. We introduce ORACLE-Grasp (O-Grasp), a novel grasping framework (Figure 1) that treats an LMM as a semantic oracle to guide grasp selection. Rather than relying on low-level image features or geometric heuristics, O-Grasp reformulates the grasping problem as an iterative, prompt-driven reasoning process, enabling human-like, task-aware grasp predictions even for novel and amorphous objects. Our approach requires no data collection, dataset curation, or pretraining, operating in a fully zero-shot manner by leveraging an existing, general-purpose LMM. Through a dual-prompt approach, O-Grasp extracts both high-level scene context and localized candidate regions, framing grasp selection as a structured, context-driven decision rather than free-form generation tasks.
The prompts are designed to be general and autonomous, allowing O-Grasp to operate without human intervention at inference time, in contrast to prior approaches that require human guidance through language commands. Critically, the use of chain-of-thought [32] reasoning initiated by the first prompt encourages the model to produce task-driven, human-like grasps that align with intuitive manipulation strategies. Although O-Grasp does not explicitly optimize for grasp stability, experimental results demonstrate that the predicted grasps are physically feasible and lead to high robot pick up success rates. Additionally, the framework integrates an early stopping mechanism to improve efficiency and a depth-based refinement step to enhance grasp reliability when depth information is available.
II Method
TABLE I: The two prompts used to provide context and locate a region to grasp
(a) Scene Context Prompt (SCP) Please provide a short, concise description of the principal object present in the image, focusing on the parts of the object that could be used for grasping. If the object has a clear handle or grip, mention that; if not, describe a cylindrical or otherwise ergonomically graspable section of the object. Avoid extraneous details unrelated to how one might physically grasp the object. Keep your answer to one sentence.
(b) Grasp Region Prompt (GRP) Based on the following image context: <SCP response>, analyze the provided image and determine the optimal grid cell, from (0,0) to (columns, rows - 1), that corresponds to the best grasping area for the object. Focus exclusively on the object (ignore all background and surrounding elements).
CONSIDER THE FOLLOWING:
   1. Prioritize areas that resemble handles or have handle-like features.
   2. If no handle is present, select the most stable area.
   3. Avoid areas that could interfere with the object’s functionality.
IMPORTANT:
   1. Your response MUST follow exactly the format below.
   2. DO NOT include any additional text, markdown formatting, or commentary.
OUTPUT FORMAT:
   GRID_CELL: <cell_number>
   EXPLANATION: <brief explanation of your choice>
While state-of-the-art LMM offer powerful semantic understanding and zero-shot object recognition, they fall short when applied to fine-grained part localization for robotic grasping. Outputs such as bounding boxes and masks often lack the spatial precision to isolate specific graspable parts, show inconsistency in the exact region they highlight across similar inputs, and do not convey the orientation required for proper gripper alignment. For instance, a handle might be detected with a bounding box that is loosely aligned, oversized, or arbitrarily rotated, making it unclear how the gripper should be positioned. Even segmentation masks, while richer in detail than boxes, often yield imprecise or fragmented regions that shift with minor input changes, offering no consistent contour or graspable axis. These limitations undermine the reliability of grasp execution, where even small errors in location or angle can lead to failure. To address this, this section introduces our method, which restructures grasp localization as an iterative decision process, enabling accurate and robust grasp prediction without explicit training.
II-A Problem Definition
We address the task of enabling a robotic arm to grasp objects placed within its workspace. The robotic arm is equipped with an RGB-D camera which can be mounted on the robot’s end-effector or somewhere in the workspace with the required calibration. We assume a single object is observed by the camera within the workspace. Although this assumption simplifies the setup, it could be relaxed in future work by incorporating context-aware reasoning based on the task or user-provided verbal instructions. The robot has no prior knowledge of the observed objects or the object to be grasped. Given an RGB image $\mathbf{I}_{i}$ and corresponding depth image $\mathbf{D}_{i}$ taken by the camera where the object is observed, the robot is required to pick up the object with a task-driven grasp. It is important to note that a standard RGB camera can be used in place of an RGB-D camera; however, this excludes the possibility of depth-based grasp refinement, which is discussed in a later section. Compensating for the lack of depth data would require additional techniques, such as visual servoing, which are beyond the scope of this work.
II-B Grasp Candidate Generation
LMMs, while proficient at generating natural language descriptions and answering image-related queries, exhibit limitations in spatial precision when embedding images into their latent space. This deficiency hinders their ability to produce pixel-accurate oriented bounding boxes for specific object parts, such as ’the handle of the mug’ or ’the handle of the drill’. Direct bounding box generation requests often yield imprecise or inconsistent outputs. Hence, we propose a novel strategy within the O-Grasp framework to guide the LMM. We first define the image tiling operation $\Gamma_{a\times b}(\mathbf{I}_{j})$ as the process of discretizing the image $\mathbf{I}_{j}$ into an $a\times b$ grid of equally sized tiles. In doing so, the problem becomes a multiple–choice question of candidate regions, where the LMM can select a tile instead of a free–form generation task. This approach is particularly effective for objects without obvious handles (e.g., a cardboard, a wallet or a shoe) and leads to more human–like grasp selections.
Figure 2: Illustration of a grasp candidate generation given an image $\mathbf{I}_{i}$ of a remote control object. Red cells on the grid denote selection for grasping by the LMM. After $m=3$ iterations with the GRP prompt over different grid configurations, stopping criterion (2) is satisfied leading to an additional $m$ iterations on a cropped image. Finally, the position $\bar{\mathbf{p}}$ and orientation angle $\theta$ of the required grasp is outputted.
Dual-prompt masking. To guide the LMM, we propose the use of a dual-prompt process that leverages tool calling capabilities. The first prompt, referred to as the Scene Context Prompt (SCP), is paired with the original image $\mathbf{I}_{i}$ and instructs the model to generate a concise description of the primary object in the scene, with an emphasis on grasp-relevant features. Here, the SCP acts as a tool call that extracts high-level semantic context, helping the model to reason more effectively about where and how to grasp the object. An example of this prompt is shown in Table I.a. This step encourages the model to engage in a chain-of-thought reasoning process about the scene, producing a contextual response that subsequently informs the subsequent prompt. The second prompt, termed Grasp Region Prompt (GRP) and seen in Table I.b, aims to approximate a grasp location. To overcome the spatial imprecision of LMMs and facilitate effective reasoning over grasp locations, we discretize the image space into a set of discrete candidate areas. It is rolled out in multiple iterations to acquire a probabilistic estimation of the grasp region. In each iteration, the GRP is paired with an image $\Gamma_{u\times v}(\mathbf{I}_{i})$ such that $u,v\in\{3,4,\dots,9\}$. The values for $v$ and $u$ are selected without repetition, starting with coarse grids and gradually refined through additional GRP queries. The GRP incorporates the contextual output of the SCP to instruct the LMM to choose the optimal tile from the discretized image (Figure 2). Repeating the GRP procedure over $K$ iterations with varying grid configurations produces a set of candidate regions =$\mathcal{B}=\{B_{1},B_{2},\dots,B_{K}\}$, where $B_{j}$ denotes a rectangular pixel mask over the image, selected by the LMM during iteration $j$. Discretizing the image space into a grid of candidate areas enables a more robust and simple selection of grasp regions.
Grasp position and orientation. For each pair of overlapping rectangular masks $B_{j}\in\mathcal{B}$ and $B_{k}\in\mathcal{B}$, we compute their Intersection-over-Union (IoU), denoted as $C_{ik}=IoU(B_{j},B_{k})$. If $C_{ik}>\gamma$, for a predefined threshold $\gamma$, the intersection region $B_{j}\cap B_{k}$ is added to the $\mathcal{B}$. A higher IoU value suggests a more promising grasp region, prompting the inclusion of an additional rectangle that reinforces areas with high grasp potential. After $K$ iterations and the inclusion of $s$ overlapping regions, the set $\mathcal{B}$ contains a total of $K+s$ rectangular pixel masks. In the subsequent step, the center coordinate $\mathbf{p}_{j}\in\mathbb{N}^{2}$ of each mask $B_{j}\in\mathcal{B}$ is computed and collected into a set $\mathcal{M}$, representing the centers of all masks. The centroid $\bar{\mathbf{p}}$ of set =$\mathcal{M}=\{\mathbf{p}_{1},\ldots,\mathbf{p}_{K+s}\}$ determines the final grasp point. Furthermore, Principal Component Analysis (PCA) is applied to the set of center points in $\mathcal{M}$ to extract the principal eigenvector $\mathbf{v}^{*}=(v^{*}_{x},v^{*}_{y})$ associated with the largest eigenvalue. The grasp orientation angle $\theta$ can be extracted from
$\theta=\operatorname{arctan2}(v^{*}_{y},v^{*}_{x}).$ (1)
Consequently, we acquire the required grasp position $\bar{\mathbf{p}}$ and orientation $\theta$ with respect to the image’s coordinate frame.
Early stopping criterion. The above process is nominally set to run for $K$ iterations. To reduce processing time and to mitigate the potential effect of multiple grasping regions, we incorporate an early stopping mechanism which also enforces focus on a specific part of the object. This mechanism assesses the spatial distribution of the rectangular mask centers as an indicator of convergence. After $m<K$ iterations, the centroid $\bar{\mathbf{p}}$ of the current set of center points ${\mathbf{p}_{1},\ldots,\mathbf{p}_{m}}$ is computed. Let $10<\rho<$ 1 be a predefined factor, and let $D$ denote the image diagonal. If the maximum distance from any center point to the centroid satisfies
$\max_{i\in\{1,\ldots,m\}}\|\mathbf{p}_{i}-\bar{\mathbf{p}}\|<\rho D,$ (2)
an additional $m$ iterations are performed on a cropped region of the image centered around the object. The cropping bounds are defined based on the largest diagonal that fits within the union of the $m$ selected grid cells ${B_{1},\ldots,B_{m}}$. If condition (2) is again satisfied after these additional iterations, the process terminates early. Otherwise, the procedure continues until all $K$ iterations are completed. This early stopping strategy not only accelerates the overall pipeline execution but also enhances robustness by ensuring that only consistent and reliable grasp candidates contribute to the final decision.
Orientation refinement. When the object’s graspable region is arbitrarily rotated relative to the image axes, the grid structure becomes misaligned, resulting in inadequate coverage and suboptimal mask fitting. This misalignment can significantly affect the accuracy of the process, particularly when the object spans a large portion of the image. To address this, we utilize PCA on the set of center points to estimate their dominant orientation. If the resulting orientation angle $\alpha$ exceeds a predefined threshold $\alpha_{min}$, the image is rotated by $-\alpha$ to better align the object with the grid structure. A new set of $m$ iterations is then executed. The resulting pixel masks are subsequently transformed back into the original image coordinate system via inverse rotation by $\alpha$. This refinement improves the estimation of both the grasp point and its orientation, enhancing the system’s precision and robustness in challenging edge cases.
II-C Grasp Position Refinement
The above process considers approximation of the required grasp point using solely RGB planar perception. Also, it uses several grid approximations that cannot be thoroughly refined due to a long computational time. Hence, the outputted grasp point may slightly deviate from the object pointing at an empty space rather than the solid structure of the object. For example, when the grasp pipeline predicts a position near the center of a kettle’s handle, which is often concave and semi-circular as illustrated in Figure 3, the depth sensor may incorrectly report an exaggerated depth value due to capturing the background through the hollow region.
To ensure a safe and accurate grasp, we employ a heuristic depth-based grasp refinement procedure that evaluates the local geometry around the predicted grasp point. Using the depth image $\mathbf{D}_{i}$, the system projects the expected closing area of the gripper based on the depth $z_{\bar{\mathbf{p}}}$ at the predicted grasp point $\bar{\mathbf{p}}$. Let $r_{ee}$ be the required clearance radius of the robot’s end-effector, and $f$ be the the camera’s focal length, the clearance radius in image units (i.e., in pixels) at depth $z$ is computed by
$r(z)=\frac{f\cdot r_{ee}}{z}.$ (3)
This radius determines the region that the gripper fingers will occupy in the image, scaled inversely with depth.
Figure 3: Illustration of a case where the outputted grasp position (circular red marker) $\bar{\mathbf{p}}$ slightly missed the kettle’s handle in the RGB image, leading to wrong depth estimation $z_{\bar{\mathbf{p}}}$. Circular regions (such as $S_{a}$ and $S_{c}$ with radii $r_{a}$ and $r_{c}$, respectively) along the line connecting the camera and original $\bar{\mathbf{p}}$ are checked for improved depth estimation toward refined grasp position $\bar{\mathbf{p}}^{*}$ (circular green marker) at depth $z^{*}$.
To verify the feasibility of the grasp, we sample $w$ depth values $\mathcal{Z}=\{z_{j}\}_{j=1}^{w}$, where each value is defined by $z_{j}=\frac{j}{w}\cdot z_{\bar{\mathbf{p}}}$. For each $z_{j}\in\mathcal{Z}$, we define a circular region $S_{j}\subset\mathbf{D}_{i}$ in the depth image $\mathbf{D}_{i}$, centered at the original grasp point $\bar{\mathbf{p}}$ and with radius $r(z_{j})$. Within each region $S_{j}$, we extract the minimal depth value $\tilde{z}_{j}\in S_{j}$, along with its corresponding pixel location $\tilde{\mathbf{p}}_{j}\in\mathbb{N}^{2}$. The final refined grasp position is then defined as $\bar{\mathbf{p}}^{*}=\tilde{\mathbf{p}}_{t}$, where
$t=\operatorname*{arg\,min}_{j\in\{1,\ldots,w\}}\tilde{z}_{j},$ (4)
and the corresponding depth is $z^{*}=z_{t}\in S_{t}$. This procedure ensures that the grasp targets the closest valid surface within the line connecting the camera and the original outputted position $\bar{\mathbf{p}}$. Hence, the robot would target a physically valid and reachable region, avoiding misleading grasp locations due to insufficient grid refinements.
III Experiments
In this section, we evaluate the performance of O-Grasp in predicting human-like grasps, and enabling robotic pick up. We adopted LLaMa 3.2-Vision (MLLaMA) [33] as the chosen LMM, with 11 billion parameters for advanced image reasoning tasks. It provides sufficient semantic reasoning power, built upon MLLaMA’s strong vision-language foundation, and its compact size allows for real-time inference on consumer-grade GPUs. Specifically, the model was deployed on a GPU cloud instance equipped with an NVIDIA H200 SXM GPU, 24 CPUs, and 377 GB of RAM. Though cloud inference is faster, preliminary tests show O-Grasp is also deployable on local hardware, with runtime as the main trade-off. Following extensive hyperparameter tuning, we set the key parameters as follows: number of GRP iterations $K=6$, IoU threshold $\gamma=0.4$, early stopping window $m=3$, and stopping factor $\rho=0.3$.
TABLE II: Position and orientation errors for O-Grasp over test-set I
Prompt variation Position Orientation
NRMSE (×$\times 10^{2}$) MAE (${}^{\circ})$
Baseline prompt 61 $\pm$ 50 -
O-Grasp w/o SCP 6.8 $\pm$ 6.1 24.3 $\pm$ 20.2
O-Grasp w/ SCP and GRP 5.1 $\pm$ 4.2 19.0 $\pm$ 14.8
Figure 4: Predicted grasps on three RGB images from test-set I, showing position (blue dot) and orientation (blue line). Objects include, from left to right: a vacuum cleaner, an AI-generated drill and a pencil. Red cells indicate regions selected by the LMM as potential grasping areas.
Figure 5: Predicted grasps (blue) versus human annotations (red) on RGB images from test-set I. Objects shown with their respective NRMSE: (a) Mug – 0.0385, (b) Saw – 0.027, (c) Adjustable wrench – 0.026, (d) Frying pan – 0.020, (e) Robot arm – 0.086, (f) Dumbbell – 0.074, (g) Hair dryer – 0.018, and (h) Suitcase – 0.046.
III-A Test-sets
[…]
III-B Evaluation metric
The proposed O-Grasp framework outputs a grasp driven by LMM reasoning. Hence, the evaluation of the predicted grasps should be compared to human reasoning, which is task-oriented. For this purpose, each image in the test-sets is annotated by an uninvolved human participant. These annotations were collected using a dedicated Graphical User Interface, allowing annotators to click on potential grasp locations and draw lines representing the intended grasp direction. The participant was instructed to subjectively annotate potential grasp centers along with the corresponding orientations, reflecting natural hand alignments for each grasp. Multiple annotations were allowed to capture different task-dependent grasp strategies - for example, a drill could be grasped by its handle for a drilling task, or by its cylindrical body when intended for handover. This process yields human annotated grasp positions $\{\bar{\mathbf{p}}_{an,1},\bar{\mathbf{p}}_{an,2},\ldots\}$ (in pixel units) and orientations $\{\theta_{an,1},\theta_{an,2},\ldots\}$ for each image in the test-sets, denoting human-like and task-oriented grasps.
The predicted grasps by O-Grasp are evaluated against human annotations. In test-set I, where no metrical data is available and object scales vary, we compute the Normalized Root Mean Squared Error (NRMSE) of the grasp position by dividing the planar distance between the predicted position $\bar{\mathbf{p}}$ and closest annotated position $\bar{\mathbf{p}}_{an,j}$ by the diameter $d$ of the object’s bounding circle: $e=\frac{|\bar{\mathbf{p}}-\bar{\mathbf{p}}_{an,j}|}{d}$. This yields a unit-less, scale-invariant measure of positional accuracy. For test-set II, which includes depth data, we report standard RMSE in metric units. In both test sets, orientation accuracy is evaluated using the Mean Absolute Error (MAE) between the predicted angle $\theta$ and annotated angle $\theta_{an,j}$ across all images.
Figure 6: Position NRMSE and computation runtime with regard to the number of GRP queries $K$ exerted in O-Grasp.
TABLE III: Grasp prediction accuracy and pick up success rates for test-set II
Object Position RMSE (mm) Orientation Success
w/o GR w/ GR MAE () Rate (%)
Coffee bag 34.8 8.9 1.01 80
Phone charger 19.91 4.0 1.51 100
USB cable 21.49 25.0 6.29 100
Brush 3.3 3.3 0.72 100
Cardboard box 29.41 8.1 3.74 80
Briefcase 30.6 21.0 8.22 60
Glue gun 35.11 7.2 15.38 80
Saw 17.5 12.3 2.6 80
Kneading knife 3.0 3.0 7.84 100
Toothpaste 8,9 8.9 5.23 100
Strengthener 15.0 18.8 18.85 80
Dustpan 30.6 11.7 7.37 80
Rubber duck 22.51 5.2 2.64 100
Stapler 10.82 9.2 13.54 100
Plastic cylinder 29.77 6.5 13.05 80
Plastic cup 27.58 16.9 10.72 100
Desk C-Clamp 18.97 19.0 4.76 80
Remote controller 19.21 2.3 4.92 100
Soldering holder 21.14 10.4 7.2 80
Bubble level tool 19.12 4.3 9.37 80
Total 20.9$\pm$9.2 10.3$\pm$6.4 7.5$\pm$5.3 88
* GR - Grasp Refinement based on depth perception (Sec. II-C)
Figure 7: Real-world objects from test-set II being grasped by a robot arm using predictions from O-Grasp. For each object, the figure shows a snapshot of the robot executing the grasp and the corresponding RGB image captured by the on-board camera. The demonstrated objects include: (a) cardboard box, (b) coffee bag, (c) six-arm soldering holder, (d) toothpaste tube, (e) dustpan, and (f) Plastic cylinder. In some cases, the initial predicted grasp (in green) was refined using depth data (final grasp shown in red), as described in Section II-C. The green circles mark the clearance region with radius $r(z_{\bar{\mathbf{p}}})$.
III-C Grasp error evaluation
We begin by evaluating grasp accuracy against human annotations in test set I. We include a baseline where the LMM is prompted to return a bounding box coordinates, with the grasp position taken as the box’s center. Though lacking orientation, this serves as a reference for NRMSE comparison. We also compare performance with and without the SCP. Without it, the model receives only the GRP, which lacks contextual cues. Table II reports the NRMSE and MAE for test set I. The baseline results indicate that directly prompting the LMM for a bounding box yields imprecise and often irrelevant regions. Note that grasp refinement was not applied to test-set I due to the absence of depth data, and the reported results exclude this step. The results demonstrate improved accuracy when the SCP is used alongside the GRP, compared to using the GRP alone. This highlights the significant impact of the SCP: by prompting the model to reason about the scene, it encourages context-aware, task-relevant grasp region selection. Figure 4 demonstrates predicted grasps over three RGB images along with an heatmap representing the LMM’s grid selections. Similarly, Figure 5 presents the grasp prediction by O-Grasp along with the human-annotated grasps. The predicted grasps reflect human-like reasoning, focusing on task-relevant regions such as handles.
We further evaluate the inference time required for generating a grasp prediction. First, the duration of a single GRP query is measured as a function of grid size $u\times v$, executing 500 GRP queries over test-set I with $u$ and $v$ sampled uniformly. The average inference time is $0.93\pm 0.07$ seconds per query. The low standard deviation suggests that GRP inference is largely insensitive to grid resolution, allowing flexible adjustment with minimal computational cost. The average total inference time for O-Grasp to predict a complete grasp over the entire test set is $7.73\pm 2.05$ seconds. This latency can be further reduced with faster GPUs or using multiple ones. Notably, 89% of O-Grasp runs terminated early due to the early stopping criterion, underscoring its practical efficiency. To illustrate the trade-off between grasp accuracy and runtime, Figure 6 shows the position NRMSE and runtime on test-set I as a function of the number of GRP queries $K$, revealing improved accuracy at the cost of increased computation.
III-D Robot grasping experiments
In the following experiment, we assess the effectiveness of O-Grasp in guiding a robotic arm to grasp real-world objects. We use a Franka Research 3 robot arm equipped with a simple parallel gripper, and a ZED-Mini camera mounted at the base of the gripper. For each grasp attempt, the arm first positioned the camera to capture an overhead image of the table. O-Grasp then processed this image to predict the grasp point and orientation, which the robot subsequently used to attempt a pick up by simply employing ROS MoveIt motion planning framework.
The evaluation covers all 20 everyday objects from test-set II, with five grasp trials per object. As before, a human annotator labeled the grasp points and orientations without observing the robot’s actual performance. Table III reports the position RMSE (with and without depth-based grasp refinement), orientation MAE, and the robot’s pick up success rate across trials. The availability of depth data in test-set II enables grasp refinement, which significantly improves positional accuracy by allowing precise adjustments before execution. Figure 7 shows snapshots from several grasping trials of the robot, displaying real-time RGB inputs alongside O-Grasp’s predicted grasps.
The results highlight O-Grasp’s robust performance across a diverse set of real-world objects. For instance, the USB cable and toothpaste tube were grasped with high success rates and showed low orientation errors, demonstrating O-Grasp’s effectiveness on both flexible and deformable items. The rubber duck, despite its irregular shape, also achieved high success rate with minimal position error. In contrast, the briefcase resulted in the lowest success rate and a relatively high orientation error. This may be attributed to its large overall size compared to the small handle, potentially requiring additional GRP iterations to identify a more focused grasp region. Similarly, the cardboard box showed the highest position error, likely due to its many potential graspable regions, which increase ambiguity in grasp selection. Despite such challenges, O-Grasp maintains high accuracy and reliability overall, with 88% average success across all grasping attempts.
IV Conclusions
This work presented O-Grasp, a hybrid grasping framework that integrates the semantic reasoning capabilities of LMMs with classical computer vision techniques to achieve robust, task-relevant grasp prediction on unknown objects. By introducing a structured dual-prompt interaction, using SCP for high level reasoning and GRP for spatial selection, we transform the inherently open-ended grasping task into a multiple-choice problem that LMMs can solve reliably. Through pre-cutting the image into candidate regions, aggregating results with IoU-based weighting and PCA, and applying depth-based clearance checks, O-Grasp effectively overcomes the spatial resolution limitations typical of LMMs. Additional mechanisms such as orientation refinement and early stopping contribute to both robustness and computational efficiency, enabling high success rates in both synthetic and real-world grasping experiments.
Despite these promising results, several limitations remain. The inference cost associated with LMMs is still substantial, posing challenges for real-time, high-frequency control loops. Future improvements could involve model compression, distillation, or leveraging more efficient architectures. Moreover, O-Grasp currently operates from a single camera viewpoint at inference time. While the model can semantically reason about objects and select plausible grasp regions, it does not actively seek alternative perspectives when critical features are occluded. Future extensions incorporating multi-view fusion or active exploration strategies could significantly enhance the system’s adaptability, bringing robotic grasping closer to human-like flexibility.